# Fast and Accurate Algorithms to Calculate Expected Modularity in Probabilistic Networks

### v2

@@ Abstract
Modularity maximization is a widely used community detection technique for deterministic networks. However, little research has been performed to develop efficient modularity calculation algorithms for probabilistic networks. Particularly, it is challenging to efficiently calculate expected modularity when all possible worlds are considered. To address this problem, we propose two algorithms, namely $\mathrm{PWP}^{\mathrm{EMOD}}$ and $\mathrm{APWP}^{\mathrm{EMOD}}$, partitioning the possible worlds based on their modularities to significantly reduce the number of probability calculations. We evaluate the accuracy and time efficiency of our algorithms through comprehensive experiments.
Keywords Modularity calculation $\cdot$ probabilistic networks $\cdot$ algorithms
1 Introduction
@@
Network data is often associated with uncertainty, which can arise because of randomness of systems under study, or inaccuracy of measurements. The first type of uncertainty can be found for example in computer networks, where network links can be unreliable. Inaccuracy of measurements is ubiquitous, for example, when estimating interaction probabilities in protein networks [1, 2, 3] or assessing the existence of social relations in social networks [4, 5]. To address these challenges, uncertain networks can be modeled as probabilistic graphs, where edges are associated with probabilities representing the likelihood of their existence [6, 7, 8].
This work is motivated by the importance of commu- nity detection, that is the identification of sets of nodes better connected to each other than to nodes outside the set, in the presence of uncertainty. As for deterministic networks, the ability to identify communities has numerous applications across various fields where a network is an appropriate data representation and uncertainty is common. For instance, in the field of biology, finding communities in probabilistic networks is crucial to predicting protein complex memberships [9]. In the context of intermittently connected networks, community detection can be used to determine delivery subgraphs [10].
@@
More specifically, this paper focuses on the important problem of calculating expected modularity in probabilistic networks. Modularity is a measure of the ratio of edges falling within groups minus approximately the expected number in an equivalent network with edges placed at random [11]. Higher values of modularity indicate better community structures. Therefore, despite some limitations of modularity as an objective function [12, 13, 14] , modularity optimization [15] has emerged as one of the most popular approaches for finding communities and core-periphery structures in deterministic networks.
@@
However, using modularity with probabilistic networks is challenging, and there are only a few studies on how to extend modularity to this context. One way of extending modularity to probabilistic networks is to calculate it over all possible worlds [16], corresponding to all the permutations of the edges in the network, and calculate its expected value. This approach is accurate but computationally expensive as it involves calculating modularity $2^{|E|}$ times, where $|E|$ is the number of edges in the network. In [3], the authors first sample $K$ graphs from all possible worlds, then calculate modularity for each sampled graph and take the average value. This approach can be very fast, but prioritizes execution time over accuracy: one can speed up the execution by choosing a lower $K$, but at the cost of obtaining a potentially inaccurate result (due to insufficient sampling) without a guaranteed approximation error. Other ways of extending modularity to probabilistic networks include regarding edge probabilities as edge weights or setting a threshold to convert probabilistic networks to deterministic networks by removing edges with probability lower than the threshold. However, these approaches have limitations. Regarding probabilities as weights generally leads to wrong results, as we show in our experiments. Setting a suitable threshold is also a complicated problem as it requires prior knowledge of the network structure and the chosen threshold greatly affects the resulting modularity [17].
In fact, in Section 4 we experimentally evaluate all the aforementioned approaches, and show that there is no known simple and general way of computing the expected value of modularity other than enumerating and computing modularity in all possible worlds, without a risk of obtaining inaccurate results.
In this paper, to avoid computing modularity in $2^{|E|}$ possible worlds, we propose an intuitive method for expected modularity computation, partitioning all possible worlds based on their modularities so that for each partition we can (1) easily compute the expected value of modularity for that partition, and (2) quickly compute or estimate the probability of that partition.
Please note that while this work is motivated by the importance of community detection, here we focus on the foundational problem of computing expected modularity given an input clustering (that is, modularity computation) and not on how to use modularity to identify good clusterings (that is, modularity optimization). In this paper a clustering is always given as an input of the algorithm, and expected modularity is computed for the input clustering.
@@
Our contribution. We use an expression for modularity calculation based on communities instead of nodes. We exploit this expression to design efficient algorithms for calculating expected modularity in probabilistic networks. Our algorithm reduces time complexity from exponential to polynomial compared to the brute-force approach. Different from e.g. sampling, whose accuracy depends on the number of samples, our method always returns an accurate result, and its execution time does not depend on the edge probability distribution. Thanks to its ability to return an accurate result in polynomial time, our method also allows us to evaluate the traditional approaches used to analyse probabilistic networks but not providing guarantees on their accuracy, that is, sampling, thresholding, and weighting, when used to estimate expected modularity. We also show that our method is robust with regard to different network structures, and study the factors influencing its running time.
@@
Roadmap. The rest of the paper is organized as follows: after reviewing the background in Section 2, we propose our algorithms in Section 3. In Section 4, we present a thorough simulation-based experimental evaluation on random networks and real-world networks with different properties. We summarize and discuss our results in Section 5.
2 Background
2.1 Probabilistic networks and relevant measures
2.1.1 Probabilistic networks
Consider a probabilistic graph $\mathcal{G}=(V,E,p)$, where $V$ corresponds to the set of nodes in $\mathcal{G}$, $E$ represents the set of edges, and $p:E\rightarrow(0,1]$ is a function that assigns probabilities to edges. We use $e_{ij}$ to indicate the edge between nodes $i$ and $j$, $p_{ij}$ as a shorthand for $p({e_{ij}})$, and we assume that $|E|=m$.
One can consider a probabilistic graph as a generative model for deterministic graphs. A deterministic graph $G$ is generated by $\mathcal{G}$ connecting two nodes $i,j$ via an edge with probability $p_{ij}$. We write $G\subseteq\mathcal{G}$ to denote that $G$ is generated by $\mathcal{G}$, and we consider the edge probabilities independent to each other [18, 19]. We assume possible worlds semantics [16], which corresponds to the set $\{G=(V,E_{G})\}_{E_{G}\subseteq E}$ of all possible deterministic graphs generated by $\mathcal{G}$ with their associated probabilities. The probability of observing such a set is:
There are $2^{m}$ distinct graphs that can be generated by $\mathcal{G}$.
2.1.2 Modularity
Broadly speaking, one solution for community detection in deterministic networks is to maximize the modularity score, which rewards partitions which capture many of the edges within communities and penalizes partitions with few or large communities. Giving a community labeling of vertices x, the modularity for an undirected network is given by:
where $M$ is the number of edges, $A$ is the adjacency matrix, $k_{i}$ represents the degree of node $i$, $x_{i}$ is the label of node $i$, and $\delta(x_{i},x_{j})$ is the Kronecker delta function, which equals 1 when its arguments are the same and 0 otherwise.
Modularity ($Q$) can also be expressed in the following form based on [15]:
2.1.3 Expected modularity
A deterministic graph $G_{w}$ generated by $\mathcal{G}$ can be considered as one possible world. From Eq. 1, we know the probability of each possible world in $\mathcal{G}$. Let Q be the discrete distribution of $Q$ on the $2^{m}$ possible worlds. The expected value of Q is:
$\begin{split}E(\textbf{Q})=\sum_{w=1}^{2^{m}}Q_{w}Pr(w),\end{split}$ (5)
where $Q_{w}$ is the modularity value of deterministic graph $G_{w}$ (that is, possible world $w$).
2.1.4 Entropy ratio
The entropy [20, 21] $H(\mathcal{G})$ of a probabilistic graph $\mathcal{G}$ is defined as the joint entropy of its edges $H(e)$ for all $e\in E$. Here we use different entropies to represent the different levels of uncertainty of a network. The larger entropy, the more uncertainty. Due to the edge independence, the formula of entropy is $H(\mathcal{G})=-\sum_{i<j}p_{ij}\log p_{ij}-\sum_{i<j}q_{ij}\log q_{ij}$, where $q_{ij}=1-p_{ij}$. To easily compare graphs with different sizes, we normalize such entropy by dividing by the number of edges. We call this entropy ratio, where entropy ratio =$\frac{|\mathcal{H}(\mathcal{G})|}{m}$, and the range is [0,1].
2.2 Existing computation methods
2.2.1 Brute-force method
One brute-force method to calculate expected modularity of probabilistic graph $\mathcal{G}$ is by dividing such probabilistic network into $2^{m}$ possible worlds [10], calculating the modularity on every possible world and their expected value according to Eq. 5. Therefore, the time complexity of this method is at least exponential on the size of the graph.
2.2.2 Sampling method
We use a sampling method called Monte Carlo (MC) estimator, which is introduced in [22]. Specifically, the MC estimator first generates $\theta$ samples from the probability distribution over the possible worlds, then calculates modularity for each sampled graph and computes the average modularity value:
$\begin{split}Q=\frac{1}{\theta}\sum_{t=1}^{\theta}Q_{t}.\end{split}$ (6)
Once the parameter $\theta$ has been specified, creating samples from the ensemble is straightforward. The detailed algorithm is depicted in Algorithm 1.
input : The number of graphs to sample $\theta$; nodes $V$; edges $E$; edges probabilities $p$
output : Graph set $Z$
1 Initialize graph set $Z$;
2 for $z\leftarrow$ 1 to $\theta$ do
3       Create an empty graph $g$ only with nodes $V$;
4       for each pair of nodes $i,j\in E$ do
5             Draw a uniformly random number $r_{ij}\in[0,1]$;
6             if $r_{ij}\leq p_{ij}$, then add the edge ($i$,$j$) to graph $g$;
7
8      add $g$ to $Z$;
9
10return $Z$;
Algorithm 1 Sampling method
2.2.3 Threshold method
Thresholding is a simple method currently used in community detection in probabilistic networks. The purpose of using threshold methods is to change probabilistic networks to deterministic networks. After setting a threshold, we remove edges whose probability is lower than the threshold and consider the others as deterministic.
3 Proposed algorithms
In this section we introduce two algorithms to compute expected modularity, exactly and approximately. The main notation introduced in this section is summarized in Table LABEL:tab:notations.
@@ 3 Proposed algorithms
probabilistic graph $\mathcal{G}$
deterministic graph $G$
@@
3.1 Possible world partitioning for expected modularity
The Possible World Partitioning for Expected Modularity ($\mathrm{PWP}^{\mathrm{EMOD}}$) algorithm is a variation of the brute-force method where instead of iterating through all possible worlds we group them into partitions, so that inside each partition all possible worlds have the same value of modularity. Using the alternative definition of modularity in Eq. 3, expected modularity can be rewritten as:
@@ 3.1 Possible world partitioning for expected modularity
where $w$ indicates a possible world, $Pr(w)$ is the probability of possible world $w$,
@@
we can partition the $2^{m}$ possible worlds so that in each partition the values of $e_{c}^{w}$, $e_{c,\bar{c}}^{w}$, and $e_{\bar{c}}^{w}$ are constant. By notating $T_{x}=\max(|e_{c}^{w}|)$, $T_{y}=\max(|e_{c,\bar{c}}^{w}|)$, and $T_{z}=\max(|e_{\bar{c}}^{w}|)$, we can write:
$\begin{split}E(\textbf{Q})=\sum_{c\in\mathcal{C}}\sum_{x=0}^{T_{x}}\sum_{y=0}^{T_{y}}\sum_{z=0}^{T_{z}}Q_{c}^{xyz}Pr(xyz)\ ,\end{split}$ (9)
where $xyz$ indicates the partition (that is, the set) of possible worlds where $|e_{c}^{w}|=x$, $|e_{c,\bar{c}}^{w}|=y$, and $|e_{\bar{c}}^{w}|=z$. When $x=y=z=0$, the corresponding graph is an empty graph and $Q_{c}^{xyz}=0$. $Q_{c}^{xyz}$ is the constant value of $Q_{c}^{w}$ in all possible worlds $w\in xyz$, that is:
@@
Notice that computing expected modularity using Eq. 9 we only need to iterate over $|\mathcal{C}|T_{x}T_{y}T_{z}$ terms instead of $2^{m}$, and for each term the value of $Q_{c}^{xyz}$ can be computed in constant time. The probability of partition $xyz$ can be expressed as a product of probabilities, because $e_{c}^{w}$, $e_{c,\bar{c}}^{w}$, and $e_{\bar{c}}^{w}$ are always disjoint sets:
$\begin{split}Pr(xyz)=Pr(|e_{c}^{w}|=x)Pr(|e_{c,\bar{c}}^{w}|=y)Pr(|e_{\bar{c}}^{w}|=z)\ .\end{split}$ (11)
These probabilities can be computed using the definition of Poisson Binomial distribution, and the probabilities on the edges in $e_{c}^{w}$, $e_{c,\bar{c}}^{w}$, and $e_{\bar{c}}^{w}$ are,
@@
$\begin{split}Pr(|e_{\bar{c}}|=z)=\sum_{A\in F_{z}}\prod_{e\in A}p(e)\prod_{e\in e_{\bar{c}}\backslash A}(1-p(e)).\end{split}$ (14)
Then the joint probability $Pr(xyz)$ is,
$\begin{split}Pr(xyz)=\prod_{\varrho\in\{x,y,z\}}\sum_{A\in F_{\varrho}}\prod_{e\in A}p(e)\prod_{e\in\delta_{\varrho}\backslash A}(1-p(e)),\end{split}$ (15)
$\delta_{\varrho}=\left\{\begin{array}[]{rcl}e_{c}&\mbox{for}&\varrho=x\\ e_{c,\bar{c}}&\mbox{for}&\varrho=y\\ e_{\bar{c}}&\mbox{for}&\varrho=z\end{array}\right.$
where $F_{\varrho}$ is the set of all subsets of $\delta_{\varrho}$ of size $\varrho$.
@@
As an example of how the partitions are defined, consider the probabilistic graph in Figure 2(a). Figure 2(b) shows a tree whose leaves enumerate all partitions $xyz$ defined by community $c$. The two branches from the top triangle represent respectively all the possible worlds where $e_{c}^{w}$ contains no edges ($x=0$) and all the possible worlds where $e_{c}^{w}$ contains one edge ($x=1$), in this case edge $l_{12}$. As we move down the tree, the circles in the second level represent all the possible worlds where $e_{c,\bar{c}}^{w}$ contains no edges ($y=0$), one edge ($y=1$, in particular $l_{23}$ or $l_{24}$), and two edges ($y=2$, in this case both $l_{23}$ and $l_{24}$). The same with the diamonds in the third level representing possible worlds with a specific number of edges in $e_{\bar{c}}^{w}$. The squares at the bottom of the figure represent all partitions. Figure 2(c) shows all possible worlds in partition 7, where $|e_{c}^{w}|=0,|e_{c,\bar{c}}^{w}|=1,|e_{\bar{c}}^{w}|=2$.
$\mathrm{PWP}^{\mathrm{EMOD}}$ is detailed in Algorithm 2. In lines 6, 8 and 10, we use Eq. 12, Eq. 13 and Eq. 14 to calculate $Pr(|e_{c}^{w}|=x)$, $Pr(|e_{c,\bar{c}}^{w}|=y)$ and $Pr(|e_{\bar{c}}^{w}|=z)$ separately.
Computational complexity: The worst-case time complexity of $\mathrm{PWP}^{\mathrm{EMOD}}$ is $O(k|E|2^{|E|})$. We can break the analysis down into two parts.
Lines 3 to 10 concern the exact calculation of probabilities for each partition. The time complexities of $Pr(x)$, $Pr(y)$, and $Pr(z)$ are $O(2^{|E|})$. From Equation 9, we can see that $x$ ranges from 1 to $T_{x}$, $y$ ranges from 1 to $T_{y}$, and $z$ ranges from 1 to $T_{z}$. $T_{x}$, $T_{y}$, and $T_{z}$ sum to $|E|$, so the overall complexity is $O(|E|2^{|E|})$.
Lines 11 to 15 involve looping over all partitions, which has a time complexity of $O(|E|^{3})$. Notice that in the worst-case scenario, the number of edges in $e_{c}$, $e_{c,\bar{c}}$, and $e_{\bar{c}}$ are similar to each other. Here the number of edges in $e_{c}$, $e_{c,\bar{c}}$, and $e_{\bar{c}}$ have the same upper bound $|E|$.
Both parts are inside the community loop on line 2, so the time complexity of the whole algorithm is $O(k(|E|2^{|E|}+|E|^{3}))=O(k|E|2^{|E|})$, where $k$ is the number of communities. Notice that the time complexity solely depends on the number of communities and edges, and remains unaffected by the assignment of edge probabilities.
(a) A probabilistic graph with a community $c$.

### v3

# On the accurate computation of expected modularity in probabilistic networks

@@ Abstract
Modularity is one of the most widely used measures for evaluating communities in networks. In probabilistic networks, where the existence of edges is uncertain and uncertainty is represented by probabilities, the expected value of modularity can be used instead. However, efficiently computing expected modularity is challenging. To address this challenge, we propose a novel and efficient technique ($\mathrm{FPWP}$) for computing the probability distribution of modularity and its expected value. In this paper, we implement and compare our method and various general approaches for expected modularity computation in probabilistic networks. These include: (1) translating probabilistic networks into deterministic ones by removing low-probability edges or treating probabilities as weights, (2) using Monte Carlo sampling to approximate expected modularity, and (3) brute-force computation. We evaluate the accuracy and time efficiency of $\mathrm{FPWP}$ through comprehensive experiments on both real-world and synthetic networks with diverse characteristics. Our results demonstrate that removing low-probability edges or treating probabilities as weights produces inaccurate results, while the convergence of the sampling method varies with the parameters of the network. Brute-force computation, though accurate, is prohibitively slow. In contrast, our method is much faster than brute-force computation, but guarantees an accurate result.
keywords:
Modularity calculation, Probabilistic networks, Algorithms
Introduction
@@
Uncertainty is an inherent property when modelling a system as a network because of randomness of the system, inaccuracy of measurements, or their interpretation. System randomness can be found, for example, in computer networks where network links can be unreliable. Inaccuracy of measurements is ubiquitous, for example when estimating interaction probabilities in protein networks [1, 2, 3] or assessing the existence of social relations in social networks [4, 5]. Uncertainty is modelled associating each edge with a probability of existence, forming a probabilistic network [6, 7, 8, 9, 10, 11, 12, 13, 14]. When the existence of all edges is certain, we talk of deterministic networks[15, 16, 9, 17, 8, 18].
@@
This paper focuses on the problem of calculating expected modularity in probabilistic networks. Modularity is a measure of the ratio of edges falling within partitions minus approximately the expected ratio in an equivalent network with edges placed at random [19]. Higher values of modularity often indicate that the input partitioning provides a good representation of the modules constituting the network. Therefore, despite some limitations of modularity as an objective function [20, 21, 22, 23], modularity optimization [24] has emerged as one of the most popular approaches for the analysis of deterministic networks.
@@
In probabilistic networks, modularity is not represented by a single value but by a distribution of modularity values arising from all combinations of edges in the network. These combinations are called possible worlds, each having a different modularity and probability in general. There are no studies proposing specific algorithms to compute the expected modularity, which means that we currently have to rely on general methods. One way of computing expected modularity in probabilistic networks is to calculate it over all possible worlds [25, 15, 9, 8], and calculate its expected value. This approach is accurate; in fact, it is the only known approach producing the correct value of expected modularity, modulo numerical approximations. However, this approach is also computationally impractical as it involves calculating modularity $2^{m}$ times, where $m$ is the number of edges in the network. Another approach is to sample $\theta$ possible worlds, then calculate modularity for each sample and take the average value[3]. This approach can be very fast, but prioritizes execution time over accuracy: one can speed up the execution by choosing a lower $\theta$, but at the cost of obtaining a potentially inaccurate result without a guaranteed approximation error. Other general ways of handling probabilistic networks, which can also be applied to modularity computation, include regarding edge probabilities as edge weights or setting a threshold to convert probabilistic networks to deterministic networks by removing edges with probability lower than the threshold. However, we show that these approaches are not appropriate for this specific task. Regarding probabilities as weights generally leads to wrong results, as we show in our experiments. Setting a suitable threshold is also a complicated problem as it requires prior knowledge of the network structure and the chosen threshold greatly affects the resulting modularity [16]. In fact, in this paper we experimentally evaluate all the aforementioned approaches, and show that there is no known simple and general way of computing the expected value of modularity other than enumerating and computing modularity in all possible worlds, without a risk of obtaining inaccurate results.
@@
Because of the limitations of existing approaches, in this paper we also introduce a new method for expected modularity computation. The novelty of this work is a new approach to exactly compute the modularity distribution without enumerating all possible worlds. Our method consists in partitioning the possible worlds so that for each partition we can (1) easily compute the expected value of modularity for the possible worlds in that partition, and (2) quickly compute the probability of the partition. The intuition behind our method is that, given a network and a clustering of its nodes, the value of modularity only depends on the number of edges inside and across communities, and not on the specific nodes incident to each edge. Therefore, we can group all possible worlds with the same numbers of in- and across-community edges into the same partition, obtaining that all possible worlds in each partition will have the same value of modularity. This would then also be the same as the expected modularity for the partition. This approach allows us to compute the correct value of expected modularity without having to compute modularity in all $2^{m}$ possible worlds, and to apply methods for the fast computation of probabilities from the Poisson Binomial distribution to the problem of expected modularity computation. Our method reduces time complexity from exponential to polynomial compared to the brute-force approach. Different from e.g. sampling, whose accuracy depends on the number of samples, our method always returns an accurate result, and its execution time does not depend on the edge probability distribution. Thanks to its ability to return an accurate result in polynomial time, our method also allows us to evaluate the traditional approaches used to analyse probabilistic networks but not providing guarantees on their accuracy, that is, sampling, thresholding, and weighting, when used to estimate expected modularity.
@@
The rest of the paper is organized as follows: after reviewing the background we present all the evaluated approaches, including the new algorithm introduced in this work. Then we present a thorough simulation-based experimental evaluation on random and real-world networks with different properties. We conclude with a summary and discussion of our results. The main notation used in the paper is summarized in Table LABEL:tab:notations.
Please note that while this work is motivated by the importance of community detection, here we focus on the foundational problem of computing expected modularity given an input clustering (that is, modularity computation) and not on how to use modularity to identify good clusterings (that is, modularity optimization). In this paper a clustering is always given as an input of the algorithm, and expected modularity is computed for the input clustering. We also note that probabilistic networks and random graph models are related but distinct concepts. A probabilistic network represents a specific real-world system, and given two nodes in that network, an edge between them either exists or not in the real world. Probabilities represent our ignorance with respect to the state of the real world.
@@
probabilistic network $\mathcal{G}$
deterministic network $G$
@@
Background
Probabilistic networks
Consider a probabilistic network $\mathcal{G}=(V,E,p)$, where $V$ corresponds to the set of nodes in $\mathcal{G}$, $E$ represents the set of edges, and $p:E\rightarrow(0,1]$ is a function that assigns probabilities to edges. We use $e_{ij}$ to indicate the edge between nodes $i$ and $j$, $p_{ij}$ as a shorthand for $p({e_{ij}})$, and we notate $|E|=m$.
Possible worlds semantics interprets a probabilistic network as a set of deterministic networks, called possible worlds, each of which associated with its probability of being observed[16, 26, 25]. That is, we represent $\mathcal{G}$ with the set $\{G=(V,E_{G})\}_{E_{G}\subseteq E}$ of all possible deterministic networks in $\mathcal{G}$ with their associated probabilities. As the edge probabilities are considered to be independent of each other[27, 28], the probability of observing such a set is:
There are $2^{m}$ distinct networks in $\mathcal{G}$.
Modularity
A common way to identify communities in deterministic networks is to maximize the modularity score, which rewards solutions capturing many of the edges within communities. Giving a community labeling of vertices x, the modularity for an undirected network is given by:
where $M$ is the number of edges, $A$ is the adjacency matrix, $k_{i}$ represents the degree of node $i$, $x_{i}$ is the label of node $i$, and $\delta(x_{i},x_{j})$ is the Kronecker delta function, which equals 1 when its arguments are the same and 0 otherwise. Modularity is a function of a network and a community assignment (that is, a partitioning of its nodes). However, it is common not to show its parameters and only write the function name ($Q$), to simplify the notation.
Modularity can also be expressed in the following form[24], which is the one used in our method:
Expected modularity
A deterministic network $G_{w}$ in $\mathcal{G}$ can be considered as one possible world. From Eq. 1, we know the probability of each possible world in $\mathcal{G}$. Let Q be the discrete distribution of $Q$ on the $2^{m}$ possible worlds. The expected value of Q is:
$\begin{split}E(\textbf{Q})=\sum_{w=1}^{2^{m}}Q_{w}Pr(w),\end{split}$ (5)
where $Q_{w}$ is the modularity value of deterministic network $G_{w}$ (that is, possible world $w$).
Entropy ratio
The entropy [29, 30] $H(\mathcal{G})$ of a probabilistic network $\mathcal{G}$ is defined as the joint entropy of its edges $H(e)$ for all $e\in E$. Here we use entropy to represent the different levels of uncertainty of a network. The larger entropy, the more uncertainty. Due to the edge independence, the formula of entropy is $H(\mathcal{G})=-\sum_{i<j}p_{ij}\log p_{ij}-\sum_{i<j}q_{ij}\log q_{ij}$, where $q_{ij}=1-p_{ij}$. To easily compare networks with different sizes, we normalize such entropy by dividing by the number of edges. We call this entropy ratio, where entropy ratio =$\frac{|\mathcal{H}(\mathcal{G})|}{m}$, and its range is [0,1].
Computation methods
Brute-force
The brute-force method to compute expected modularity directly uses Eq. 5, calculating the modularity on every possible world and their expected value[31]. Therefore, the time complexity of this method is at least exponential on the size of the network. While impractical, this method returns the correct value of expected modularity, and can thus be used on small networks to evaluate the accuracy of other approaches.
Sampling
The sampling method called Monte Carlo estimator[32] first generates $\theta$ samples from the probability distribution over the possible worlds, then calculates modularity for each sample $t$ (which we notate $Q_{t}$) and computes the average modularity value:
$\begin{split}Q=\frac{1}{\theta}\sum_{t=1}^{\theta}Q_{t}\ .\end{split}$ (6)
Once the parameter $\theta$ has been specified, creating samples from the ensemble is straightforward. The detailed algorithm is depicted in Algorithm 1. Note that we do not use Naive Monte Carlo sampling, that is known to have a higher variance [33].
input : The number of networks to sample $\theta$; nodes $V$; edges $E$; edges probabilities $p$
output : Set of networks $Z$
1 Initialize $Z$;
2 for $z\leftarrow$ 1 to $\theta$ do
3       Create an empty network $g$ only with nodes $V$;
4       for each pair of nodes $i,j\in E$ do
5             Draw a uniformly random number $r_{ij}\in[0,1]$;
6             if $r_{ij}\leq p_{ij}$, then add the edge ($i$,$j$) to network $g$;
7
8      add $g$ to $Z$;
9
10return $Z$;
Algorithm 1 Sampling method
Thresholding
Thresholding is a simple and general method used in community detection in probabilistic networks. Its purpose is to transform probabilistic networks into deterministic networks, so that existing methods for deterministic networks can be applied. After setting a threshold, we remove edges whose probability is lower than the threshold and consider the others as deterministic.
Possible-World Partitioning (PWP)
The Possible-World Partitioning for Expected Modularity ($\mathrm{PWP}$) algorithm, that we introduce in this work, groups the possible worlds into partitions so that all possible worlds inside the same partition have the same value of modularity. For a given community assignment $\mathcal{C}$ and using the alternative definition of modularity in Eq. 3, expected modularity can be rewritten as:
@@ 5             Draw a uniformly random number $r_{ij}\in[0,1]$;
where $w$ indicates a possible world corresponding to deterministic network $G_{w}$, $Pr(w)$ is the probability of this possible world,
@@
we can partition the $2^{m}$ possible worlds in a way that allows us to process all the possible worlds in each partition without iterating over them. In particular, the key idea of our approach to reduce computational complexity is to define these partitions so that the possible worlds within the same partition have the same modularity. We do this by defining partitions whose possible worlds have the same number of edges in $e_{c}^{w}$, $e_{c,\bar{c}}^{w}$, and $e_{\bar{c}}^{w}$, respectively. Note that these sets are disjoint and in general have different sizes. We notate $d^{xyz}$ the partition containing all possible worlds $w$ where $|e_{c}^{w}|=x$, $|e_{c,\bar{c}}^{w}|=y$, and $|e_{\bar{c}}^{w}|=z$. This partition contains $\binom{T_{x}}{x}\cdot\binom{T_{y}}{y}\cdot\binom{T_{z}}{z}$ possible worlds, where $T_{x}=|e_{c}|$, $T_{y}=|e_{c,\bar{c}}|$, and $T_{z}=|e_{\bar{c}}|$ are the (maximum) number of edges in the three parts, respectively. We can then write:
$\begin{split}{{\color[rgb]{0,0,0}E(\textbf{Q})=\sum_{c\in\mathcal{C}}\sum_{x=0}^{T_{x}}\sum_{y=0}^{T_{y}}\sum_{z=0}^{T_{z}}Q_{c}^{xyz}Pr(d^{xyz}}})\ ,\end{split}$ (9)
where $Q_{c}^{xyz}$ is the constant value of $Q_{c}^{w}$ in all possible worlds $w\in d^{xyz}$, that is:
@@
Notice that computing expected modularity using Eq. 9 we only need to iterate over $|\mathcal{C}|T_{x}T_{y}T_{z}$ terms instead of $2^{m}$, and for each term the value of $Q_{c}^{xyz}$ can be computed in constant time. The probability of partition $d^{xyz}$ can be expressed as a product of probabilities, because $e_{c}^{w}$, $e_{c,\bar{c}}^{w}$, and $e_{\bar{c}}^{w}$ are always disjoint sets:
$\begin{split}{{\color[rgb]{0,0,0}Pr(d^{xyz})=Pr(|e_{c}^{w}|=x\land|e_{c,\bar{c}}^{w}|=y\land|e_{\bar{c}}^{w}|=z)=Pr(|e_{c}^{w}|=x)\cdot Pr(|e_{c,\bar{c}}^{w}|=y)\cdot Pr(|e_{\bar{c}}^{w}|=z)\ .}}\end{split}$ (11)
These probabilities can be computed using the definition of Poisson Binomial distribution. In the following equations, $A$ is an element of $F_{x}$, $F_{y}$ or $F_{z}$, where $F_{x}$ is the set of all subsets of $e_{c}$ of size $x$, $F_{y}$ is the set of all subsets of $e_{c,\bar{c}}$ of size $y$, $F_{z}$ is the set of all subsets of $e_{\bar{c}}$ of size $z$.
@@
$\begin{split}Pr(|e^{w}_{\bar{c}}|=z)=\sum_{A\in F_{z}}\prod_{e\in A}p(e)\prod_{e\in e_{\bar{c}}\backslash A}(1-p(e))\ .\end{split}$ (14)
@@
As an example of how the partitions are defined, consider the probabilistic network in Figure 2. Figure 3 shows a tree whose leaves enumerate all partitions $d^{xyz}$ defined by community $c$. The two branches from the top triangle represent respectively all the possible worlds where $e_{c}^{w}$ contains no edges ($x=0$) and all the possible worlds where $e_{c}^{w}$ contains one edge ($x=1$), in this case edge $l_{12}$. As we move down the tree, the circles in the second level represent all the possible worlds where $e_{c,\bar{c}}^{w}$ contains no edges ($y=0$), one edge ($y=1$, in particular $l_{23}$ or $l_{24}$), and two edges ($y=2$, in this case both $l_{23}$ and $l_{24}$). The same with the diamonds in the third level representing possible worlds with a specific number of edges in $e_{\bar{c}}^{w}$. The squares at the bottom of the figure represent all partitions. Figure 4 shows all possible worlds in partition 7, where $|e_{c}^{w}|=0,|e_{c,\bar{c}}^{w}|=1,|e_{\bar{c}}^{w}|=2$.
$\mathrm{PWP}$ is detailed in Algorithm 2. In lines 6, 8 and 10, we use Eq. 12, Eq. 13 and Eq. 14 to calculate $Pr(|e_{c}^{w}|=x)$, $Pr(|e_{c,\bar{c}}^{w}|=y)$ and $Pr(|e_{\bar{c}}^{w}|=z)$ separately.
While $\mathrm{PWP}$ can already be considered a usable algorithm, because it reduces the number of modularity computations from exponential to polynomial, its worst-case time complexity is still $O(km2^{m})$. We can break the analysis down into two parts. Lines 3 to 10 concern the exact calculation of probabilities for each partition. The time complexities of $Pr(x)$, $Pr(y)$, and $Pr(z)$ are $O(2^{m})$. From Equation 9, we can see that $x$ ranges from 1 to $T_{x}$, $y$ ranges from 1 to $T_{y}$, and $z$ ranges from 1 to $T_{z}$. $T_{x}$, $T_{y}$, and $T_{z}$ sum to $m$, so the overall complexity is $O(m2^{m})$. Lines 11 to 15 involve looping over all partitions, which has a time complexity of $O(m^{3})$. Notice that in the worst-case scenario, the number of edges in $e_{c}$, $e_{c,\bar{c}}$, and $e_{\bar{c}}$ are similar to each other. Here the number of edges in $e_{c}$, $e_{c,\bar{c}}$, and $e_{\bar{c}}$ have the same upper bound $m$. Both parts are inside the community loop on line 2, so the time complexity of the whole algorithm is $O(k(m2^{m}+m^{3}))=O(km2^{m})$, where $k$ is the number of communities. Notice that the time complexity solely depends on the number of communities and edges, and remains unaffected by the assignment of edge probabilities.
Figure 2: A probabilistic network with a community $c$.